# SCEP Collaboration The projected sensitivity of SCEP experiment to Magnetic Monopole

### v2

@@ Abstract
The investigation of beyond-Standard-Model particles is a compelling direction in the pursuit of new physics. One such hypothetical particle, the magnetic monopole, has garnered considerable attention due to its strong theoretical motivation and potential to unveil profound physical phenomena. The magnetic monopole is intricately linked to the long-standing enigma surrounding the quantization of electric charge. In this manuscript, we propose a novel detection scenario for magnetic monopoles by employing a coincidence measurement technique that combines a room-temperature magnetometer with plastic scintillators. This setup allows for the collection of both the induction and scintillation signals generated by the passage of a monopole. The estimation of the sensitivity using a simple benchmark setup is given.
@@ I Introduction
The relentless pursuit of uncovering new physics occupies a prominent position in modern scientific exploration. Despite the remarkable success of the Standard Model (SM) of particle physics in elucidating the behavior of fundamental particles and their interactions, it is widely acknowledged that SM remains incomplete. Numerous enigmatic phenomena persist as tantalizing mysteries, including the elusive nature of dark matter, the perplexing origin of matter-antimatter asymmetry, and the unification of fundamental forces. Consequently, physicists actively engage in tireless searches for novel physics beyond the SM. This endeavor encompasses both theoretical advancements and experimental undertakings, propelling the boundaries of human comprehension and challenging existing scientific paradigms. Among the directions pursued in these explorations, the search for beyond-Standard-Model particles plays a pivotal role, compelling researchers to employ state-of-the-art detector techniques to scrutinize hypothetical particles that hold the potential to unveil the secrets of new physics. Magnetic monopole is one prominent candidates for beyond-Standard-Model particles that have garnered considerable attention within the scientific community.
@@
A magnetic monopole (MM) is a theoretical particle postulated to exist as an isolated source of a singular magnetic charge, analogous to the individual positive or negative electric charges observed in particles. Proposed by Paul Dirac in 1931 [dirac1931quantised] as a consequence of his pioneering work on the quantization of electric charge, MMs hold significance in fundamental physics as they provide a means to unify electromagnetism and explain the quantization of charge. The concept of MMs finds natural incorporation within the framework of Grand Unified Theories (GUTs) [georgi1974unity], which aim to unify the electromagnetic, weak, and strong nuclear forces. The quantization of electric charge is also explained in the framework of GUT. The observed lack of GUT-MMs in the Universe today is one of the key motivations behind the proposal of the cosmological inflation [zeldovich1978concentration, guth1981inflationary]. The search for MMs persists through various experimental approaches, prominently including ultra-low background experiments and superconducting coil-based experiments, which strive to detect the elusive presence of these MMs and further our understanding of the fundamental laws. Ultra-low background experiments are typically conducted in underground environments with kilometers of rock overburdens, providing shielding against cosmic rays. These experiments aim to detect the ionization or scintillation signals produced by MMs as they traverse the target material of the detector. Notably, the MACRO experiment [macro2002final], based on liquid scintillator technology, and the neutrino telescope IceCube [abbasi2013search] have yielded the most sensitive searches for MMs with speeds greater than 4×10-5 times the speed of light. While the ionization density caused by MMs is predicted to be significantly higher than that of background particles commonly observed in terrestrial detectors, such as muons and electrons, it is important to consider the possibility of alternative exotic particles, such as superheavy dark matter [bertone2005particle, chung1998superheavy], which could also contribute to high-density ionization. Conversely, superconducting coil-based experiments [eberhard1975improvements, cabrera1982first] focus on detecting the smoking-gun induction signals generated by MMs. However, these experiments face limitations in terms of size due to the requirement of maintaining superconducting temperatures.
This article presents a comprehensive illustration of the SCEP (Search for Cosmic Exotic Particles) experiment, with a specific emphasis on the detection perspective of MMs. We propose a novel approach utilizing a coincidence measurement technique that combines room-temperature magnetometers with plastic scintillators (PS). Such searches for GUT-MMs can be carried out at both the sea-level and even high-altitude environments, taking advantage of the significant reduction in the particle background achieved through the coincidence requirement. Moreover, carrying out such searches in the deep-space environment, such as on the moon, offers additional advantages. The lunar environment has low magnetic noise and no atmosphere, making it an ideal setting to detect non-GUT-MMs, which may have considerably lower masses that would be inaccessible to terrestrial detectors due to the Earth’s atmosphere.
@@
The fundamental concept of the detector system is illustrated in Section II. To assess the capabilities of the proposed system, we have developed a sophisticated simulation framework, which is described in Section III. The validation of the simulation framework is performed and described in Section IV. Furthermore, the anticipated background and sensitivity of the SCEP experiment to MMs are presented in Section V.
@@ II Detector Concept
A single module of the SCEP detector encompasses dedicated detection systems for both the scintillation and induction signals, as illustrated in Fig. 1. The scintillation signals are captured by PSs positioned at the top and bottom of the module. In the preliminary design, each PS module is constructed using the designs similar to the ones utilized in previous works [zhou2020time]. To guide the scintillation light, wavelength-shifting fibers are strategically incorporated within the PS module. These fibers serve the purpose of directing the emitted light to Silicon photomultipliers (SiPMs) coupled at the ends of the fibers. A preliminary simulation using the GEANT4 toolkit [agostinelli2003geant4] has been conducted to evaluate the performance of the PS module. The obtained results indicate a light yield of approximately 22 photoelectrons (PE) per MeV, thereby enabling an energy resolution of about 8.6% and 2.5% for muons at 8 MeV and Dirac MMs at 100 MeV, respectively.
![Refer to caption](x1.png)
![Refer to caption](x2.png)
@@
The induction signals resulting from the passing through of a MM are collected using a specialized apparatus that integrates an induction coil, a Helmholtz coil, and a magnetometer. The micro-current induced by the MM passing through the induction coil is subsequently directed to the Helmholtz coil, leading to the generation of an alternating magnetic field at the center of the Helmholtz coil. Subsequently, the alternating magnetic field is detected utilizing a magnetometer renowned for its exceptional sensitivity [jiang2020interference, jiang2021search]. This magnetometer is meticulously designed to exhibit an extraordinary level of sensitivity towards the variations in magnetic fields. The target material of the magnetometer is confined in a transparent gas chamber, and is polarized by the static magnetic field aligned along the axis with the assistance of a beam of bump laser. The presence of an alternating magnetic field in the plane can impact the precession of the atoms within the gas chamber. This effect manifests as the variations in the light density of a laser beam which pass through the gas chamber. More details of the magnetometer are given in Ref. [jiang2021search]. A preliminary prototype of the magnetometer can reach a detection sensitivity of 1 fT/ for the alternating magnetic field [kominis2003subfemtotesla, hong2024femtotesla]. Besides, an alternative readout scenario is being considered for the search of high-speed MM. This scenario involves a direct connection between the induction coil, an operational amplifier (OPA), and an analog-to-digital converter (ADC). Although this setup exhibits higher intrinsic noise levels compared to the use of a magnetometer as the readout method, it provides the benefits of quicker response times, compact size conducive to integration, and more cost-effective, lightweight systems.
@@ III Simulation of Signal
The signal-to-noise ratio (SNR) to Dirac MM stands as a critical parameter governing the quality of the induction signal. The SNR in this work is defined as the maximum signal amplitude squared divided by the mean-squared noise amplitude :
Larger values of SNR are preferred for higher noise rejection power.
This criterion allows for the utilization of a lower number of coil coincidences in detector array while still attaining a relatively high level of noise rejection. The SNR of the system is related to various factors, including the electrical parameters encompassing the circuitry from the induction coil to the Helmholtz coil, the prevailing temperature conditions, the signal response characteristics of the magnetometer, and other relevant factors. These factors collectively contribute to the overall SNR, influencing the system’s ability to discern and extract the desired signal amidst the presence of noise. To estimate the SNRs for MMs with various velocities and to optimize the design of the induction system, a comprehensive simulation framework has been developed which is described briefly in the following subsections. It should be noted that in this manuscript, the magnetic charge of Dirac MM is employed as a reference standard for the calculation of scintillation and induction signals attributed to a MM.
@@ III.1 Induction
The induction voltage on the induction coil is calculated assuming that the thickness of coil brings negligible effect. When a MM with velocity of passing through the induction coil with radius of , the induction voltage can be written as in Eq. 2. Eq. 2 is based on the assumption that the time is 0 when the MM reaches the coil plane (). is the transverse distance to the coil center when the MM reaches the coil plane. and represent the polar angle and azimuth angle, respectively, of the incoming MM’s direction under the spherical coordinates with the axis perpendicular to the coil plane. The Wb is the magnetic charge of Dirac MM [dirac1931quantised], and is the coil turn number. The and functions are the complete elliptic integrals of the first and second kinds, respectively. The induction signal is at maximal when the MM passes through the coil center with . The amplitude and spectral shape of the induction signals are predominantly influenced by the MM speed, the polar angle, and the transverse distance to the coil center. These dependencies are visually depicted in Fig. 2.
The interaction of charged SM particles or SM particles with magnetic moments with the induction coils can potentially result in induction signals. However, there are significant distinctions in the amplitude and spectral shape of these signals compared to those induced by the MMs. More importantly, the induction signals generated by SM particles have a vanishing time integral due to their nature as, at most, magnetic dipoles. On the other hand, common background SM particles, such as muons, neutrons, and protons, typically exhibit relativistic speeds, leading to rapid resonant induction on the timescale of approximately 10 picoseconds for a coil with a 12-centimeter diameter. The quick oscillation of voltage cannot be effectively shaped by subsequent relatively slow circuitry or reliably detected by read-out devices. Considering these factors, the induction caused by SM particles is considered to be negligible in practical scenarios.
@@ III.2 Signal shaping
![Refer to caption](x3.png)
The induction coil and the Helmholtz coil are connected in series111The induction coil is connected directly to the amplifier and ADC, but the simulation procedures are mostly identifcal.. These coils possess non-trivial resistances, capacitances, and inductances, which affect both the amplitude and temporal characteristics of the electric current within the circuit. In the signal simulation, it is assumed that the coil can be approximated as a series combination of a resistor and an inductor, paralleled by a capacitor. The circuit diagram of the induction and Helmholtz coils in the magnetometer-readout scenario, as well as of the direct read-out scenario using the ADC, is depicted in Fig. 3. In the circuit diagram, (), (), and () are the effective inductance, resistance, and capacitance, respectively, of the induction (Helmholtz) coil. represents other parallel capacitive components in the circuit, mainly the distributed capacitance of the cable and the input capacitance of the OPA. is the induction voltage, and is the induction current on the Helmholtz coil which is directly related the strength of magnetic field that is eventually captured by the magnetometer in magnetometer-readout scenario. In the alternative ADC-readout scenario, represents the voltage detected by the ADC, while = represents the electric current flowing into the ADC. is the coupling resistance of the ADC.
![Refer to caption](x4.png)
Coil | Wire type | Wire diameter [mm] | Minimal coil radius [cm] | Maximal coil radius [cm] | Turn number | Optimized SNR0 | |
@@
In the context of signal simulation framework, the electric current is determined by applying a circuit response function to the induction voltage . The Fourier transform of the electric current, denoted as , can be expressed as:
where the complex is the Fourier transform of the induction voltage. The circuit response function is analytically derived based on the effective circuit models shown in Fig. 3. A response function for a 6-cm-radius coil with 4320 turns is presented in Fig. 4. The resonant frequencies of the two readout scenarios exhibit variations owing to disparities in the circuit configurations. In particular, the inductance and capacitance of the Helmholtz coil in the magnetometer readout scenario contribute to a higher resonant frequency compared to the alternative ADC readout scenario. Among the various electric parameters, the resistance of the induction coil is identified as the most dominant factor. The resistance depends on the signal frequency , mainly due to the presence of the skin effect and the proximity effect [dowell1966effects]. However, the exact relation between the coil resistance and signal frequency cannot be analytically given due to the complexity of the coil structure. To investigate the frequency dependence of the coil resistance, in-situ measurements are conducted using an HIOKI IM3533-01 LCR meter. The LCR meter is connected to both ends of the coil to test the coil’s complex impedance at different frequencies. The magnitude is denoted as , and the phase angle is denoted as . The and have correlation with the inductance , capacitance , and resistance of the coil, which can be expressed as:
@@
where is the direct resistance of the coil, which is independent of signal frequency. The parameters and are empirical model parameters. Once the complex impedance of the coil is measured including and , a Nelder-Mead fitting algorithm is utilized to derive the resistance of the coil. In the benchmark tests, three induction coils with different radius and turn numbers are manufactured and tested. Their geometrical parameters are given in Table 1. The measured results for these induction coils are shown in Fig. 5. The waveforms of the induction electric current on the Helmholtz coil in magnetometer readout scenario and of electric current flowing into ADC in alternative readout scenario (the current in Fig. 3) are shown in Fig. 6, assuming the MM perpendicularly pass the induction coil center.
![Refer to caption](x5.png)
![Refer to caption](x6.png)

### v3

# SCEP Collaboration SCEP: a Cosmic Magnetic Monopole Search Experiment

@@ Abstract
Magnetic monopole is a well-motivated class of beyond-Standard-Model particles that could provide insights into the long-standing puzzle of the quantization of electric charge. These hypothetical particles are likely to be super heavy (1015 GeV) and be produced in the very early stages of the Universe’s evolution. We propose a novel detection scenario for the search of such cosmic magnetic monopoles, utilizing a hybrid approach that combines radio-frequency atomic magnetometers and plastic scintillators. Such setup allows for the collection of both the induction and scintillation signals generated by the passage of a magnetic monopole, which provides acceptance to the magnetic monopoles with their velocities larger than about 10-6 light speed (assuming a signal-to-noise ratio of 4) and their masses larger than approximately 107 GeV (at ). The proposed detector design has the potential to scale up to large area, enabling the exploration of the parameter space of the cosmic magnetic monopole beyond the current experimental and astrophysical constraints. It is estimated that such detector can reach current most stringent limits of the flux set by previous searches, with a signal-to-noise ratio of the induction signal larger than about 4.5, assuming an effective exposure being 20000 yearm2 and coil layer of 3.
@@ I Introduction
A magnetic monopole (MM), proposed by Dirac in 1931 [dirac1931quantised], is a theoretical particle postulated to exist as an isolated source of a singular magnetic charge, analogous to the electric charge. The MMs hold significance in fundamental physics as they provide a means to explain the quantization of electric charge. Dirac derived that the minimal magnetic charge of MM is Wb. The MM with this unit magnetic charge (denoted as Dirac MM) is the benchmark for most of the MM searches. The concept of the MMs also finds natural incorporation within the framework of Grand Unified Theories (GUTs) [georgi1974unity], which aim to unify the electromagnetic, weak, and strong nuclear forces. The quantization of electric charge is also explained in the framework of GUT. Most GUT models generally predict that large numbers of the MMs are created when the strong force decouples from the electro-weak force. The typical energy scale is 1015 GeV which also determines the mass scale of the GUT-MMs. The observed lack of the GUT-MMs in the universe today is one of the key motivations behind the proposal of the cosmological inflation [zeldovich1978concentration, guth1981inflationary].
@@
The searches for these cosmic MMs predicted by the GUT models persist through various experimental approaches, including the ultra-low background experiments and the superconducting coil-based experiments. More information about the searches of the MMs can be found in Ref. [preskill1984magnetic, patrizii2015status]. The ultra-low background experiments are typically conducted in the underground environments with kilometers of rock overburdens, providing a shielding against the cosmic-ray backgrounds. These experiments aim to detect the ionization or scintillation signals produced by the MMs as they traverse the target material of the detector. Notably, the MACRO experiment [macro2002final] and the neutrino telescope IceCube [abbasi2013search] have yielded the most sensitive searches for the MMs with speeds greater than about 4×10-5 times the light speed and with relativistic speeds, respectively. While the ionization density caused by the MMs is predicted to be higher (particulary for high-speed MMs) than that of the background particles commonly observed in the terrestrial detectors, such as muons and electrons, due to the unit magnetic charge being much larger than the unit electric one. It is important to consider the possibility of alternative exotic particles, such as the superheavy dark matter [bertone2005particle, chung1998superheavy], which could also create high-density ionization. Conversely, superconducting coil-based experiments [eberhard1975improvements, cabrera1982first] focus on detecting the smoking-gun induction signals generated by the MMs. However, these experiments face limitations in terms of their sizes due to the requirements of maintaining superconducting temperatures. It is also worth noting that MMs predicted by non-GUT models can have distinct masses, velocities, and magnetic charges (see Ref. [preskill1984magnetic, rossi1982exact, weinberg2007magnetic, mavromatos2020magnetic] for theoretical reviews). Dirac MMs with magnetic charge of is usually set as primary search target to be a conservative approach. Numerous scientific experiments have been conducted to search for such non-GUT MMs. For instance, experiments performed near particle colliders have been utilized to investigate the potential existence of the low-mass MMs [acharya2014physics, bertani1990search, abbiendi2008search, kinoshita1992search, pinfold1993search]. Additionally, experiments that collect and analyze the terrestrial and extraterrestrial samples [price1988limits, ghosh1990supermassive, bendtz2013search] have been employed to search for the MMs that may be bound to matter.
This article presents a comprehensive illustration of the SCEP (Search for Cosmic Exotic Particles) experiment, with a specific emphasis on the detection perspective of mainly the cosmic MMs. We propose a novel approach utilizing a coincidence measurement technique that uses the radio-frequency atomic magnetometers and the plastic scintillators (PSs). The primary components of this MM detector are room-temperature induction coils and PSs, which make it a cost-effective and scalable system. Simulations have shown that this detection setup is sensitive to MMs with velocities (assuming a signal-to-noise ratio of about 4) 111Further increasing the signal-to-noise ratio of the induction detection can extend the search to lower velocity region.. Such searches for the cosmic MMs can be carried out at the sea-level or high-altitude environment, rather than in an underground laboratory, taking advantage of the reduction in the charged particle background achieved through the coincidence requirement between the induction coils and PS detectors. A surface or high-altitude detector deployment can lower the mass threshold of the cosmic MM search from about 1010 GeV to 107 GeV [patrizii2019searches] (at ) compared to an underground detector, due to the reduced overburden. This will be detailed also in Sub-section V.5. Based on these advantages, this detection scheme holds potential for the large-scale deployment, surpassing the current best flux limits [macro2002final, abbasi2013search, giacomelli1984magnetic] and the astrophysical constraints [adams1993extension, lewis2000protogalactic] for the cosmic MMs.
@@
The fundamental concept of the detector system is illustrated in Section II. For estimating the sensitivities to the cosmic MMs using the proposed system, we have developed a simulation framework of the induction signal, which is described in Section III. The validation of the simulation is performed and described in Section IV. Furthermore, the estimation of the background and sensitivity of the SCEP experiment to the cosmic MMs are presented in Section V.
@@ II Detector Concept
A single module of the SCEP detector encompasses dedicated the detection systems for both the scintillation and induction signals, as illustrated in Fig. 1. The scintillation signals are captured by the PSs positioned at the top and bottom of the module. In the preliminary design, each PS module is constructed using the designs similar to the ones utilized in [zhou2020time]. To guide the scintillation light, wavelength-shifting fibers are incorporated within the PS module. These fibers serve the purpose of directing the emitted light to Silicon photomultipliers (SiPMs) coupled at the ends of the fibers. A preliminary simulation using the GEANT4 toolkit [agostinelli2003geant4] has been conducted to evaluate the performance of the PS module. The obtained results indicate a light yield of approximately 22 photoelectrons (PE) per MeV, thereby enabling an energy resolution of about 8.6% and 2.5% for muons at 8 MeV and Dirac MMs at 100 MeV, respectively.
@@
The induction signals resulted from the passage of a MM are collected using an apparatus that integrates an induction coil, a Helmholtz coil, and a magnetometer. The micro-current induced by the MM passing through the induction coil is subsequently directed to the Helmholtz coil, leading to the generation of an alternating magnetic field at the center of the Helmholtz coil, which is wound around the core gas chamber of a radio-frequency atomic magnetometer. Subsequently, the alternating magnetic field is detected utilizing this kind of magnetometer renowned for its exceptional sensitivity to magnetic field. In our detector setup, a magnetometer with a sensitivity of several fT/ to the magnetic field will suffice since the dominant noise comes from the thermal noise of the induction coil. The target material (for example, rubidium 87) of the magnetometer is confined in a transparent gas chamber, and is heated to maintain a temperature of around 200 degrees Celsius. The material is also polarized by a static magnetic field aligned along the axis with the assistance of a beam of pump laser. The presence of an alternating magnetic field in the plane can impact the precession of the atoms within the gas chamber. This effect manifests as the variations in the measured light intensity of a laser beam which pass through the gas chamber. More details of the working principles of the magnetometer are given in Ref. [lee2006subfemtotesla, savukov2005tunable, savukov2007detection, jiang2020interference, jiang2021search]. A preliminary prototype of the magnetometer can reach a detection sensitivity of 10 fT/ for the alternating magnetic field [kominis2003subfemtotesla, hong2024femtotesla]. The frequency range of the magnetometer readout is approximately from several hundred Hz to several tens of kHz, limited by the Larmor precession frequency of the target atoms in the magnetometer under the static polarization magnetic field that can be practically applied. Additionally, we consider an alternative readout configuration. In this configuration, the induction coil is directly connected to an operational amplifier (OPA) and an analog-to-digital converter (ADC). Although this configuration has higher intrinsic back-end noise levels compared to a magnetometer readout, it can reach a higher frequency range up to the MHz level, potential extending the induction sensitivity for the fast MMs. This configuration also has other advantages, such as a more compact size conducive to integration, a more cost-effective design, and reduced overall weight.
@@ III Simulation of Signal
The signal-to-noise ratio (SNR) to a single Dirac MM stands as a critical measure of the quality for the induction signal in this work. The SNR in this work is defined as the maximum signal amplitude squared divided by the mean-squared noise amplitude :
Larger values of SNR are preferred for higher noise rejection power. The SNR of the system is related to various factors, most dominantly the resistance of the induction coil, which introduces significant thermal noise. Additionally, the prevailing temperature conditions, as well as the signal response characteristics of the circuit and the magnetometer, also plays a role. To estimate the SNRs for MMs with various velocities, a comprehensive simulation framework has been developed which is described briefly in the following subsections.
@@ III.1 Induction
The induction voltage on the induction coil is calculated assuming that the thickness of the coil brings negligible effect. When a MM with the velocity of passing through the induction coil which has a radius of , the induction voltage can be written as in Eq. 2. Eq. 2 is based on the assumption that the time is 0 when the MM crosses the coil plane (). is the transverse distance to the coil center when the MM reaches the coil plane. and represent the polar and azimuth angles, respectively, of the incoming MM’s direction under the spherical coordinates with the axis perpendicular to the coil plane. is the coil turn number, the and functions are the complete elliptic integrals of the first and second kinds. The induction signal is at maximal when the MM passes through the coil center with . The amplitude and spectral shape of the induction signals are predominantly influenced by the MM speed, the polar angle, and the transverse distance to the coil center. These dependencies are shown in Fig. 2.
The interaction of charged Standard Model (SM) particles, or SM particles with magnetic moments, with the induction coils can potentially result in induction signals. However, there are significant distinctions in the amplitude and spectral shape of these induction signals compared to those induced by the MMs. More importantly, the induction signals generated by the SM particles have a vanishing time integral due to their nature as, at most, magnetic dipoles. In addition, the common background SM particles, such as the muons, neutrons, and protons, typically exhibit relativistic speeds, leading to the rapid resonant induction on the timescale of approximately 10 picoseconds for a coil with a 12-centimeter diameter. The quick oscillation of the voltage cannot be effectively shaped by the subsequent relatively “slow” circuitry and reliably detected by the read-out devices. Considering these factors, the induction caused by SM particles is considered to be negligible.
@@ III.2 Signal shaping
The induction coils possess non-trivial resistances, capacitances, and inductances, which affect both the amplitude and temporal characteristics of the electric current within the circuit. In the signal simulation, it is assumed that a single induction coil can be approximated as a series combination of a resistor and an inductor, paralleled by a capacitor. The circuit diagram of the induction and Helmholtz coils in the magnetometer-readout scenario, as well as of the direct readout scenario using the ADC, is depicted in Fig. 3. In the circuit diagram, (), (), and () are the effective inductance, resistance, and capacitance, respectively, of the induction (Helmholtz) coil. represents other parallel capacitive components in the circuit, mainly the distributed capacitance of the cable and the input capacitance of the OPA. is the induction voltage, and is the induction current on the Helmholtz coil which is directly related the strength of magnetic field that is eventually captured by the magnetometer in the magnetometer-readout scenario. In the alternative direct ADC-readout scenario, represents the voltage detected by the ADC, while = represents the electric current flowing into the ADC. is the coupling resistance of the ADC.
| Coil | Wire type | Wire diameter [mm] | Minimal coil radius [cm] | Maximal coil radius [cm] | Turn number | SNR0 | |
@@
In the signal simulation, the electric current is determined by applying a circuit response function to the induction voltage . The Fourier transform of the electric current, denoted as , can be expressed as:
where the complex is the Fourier transform of the induction voltage. The circuit response function is analytically derived based on the effective circuit models shown in Fig. 3. A response function for a 6-cm-radius coil with 4320 turns is presented in Fig. 4. The most sensitive frequency range (frequencies around the resonant peaks in Fig. 4) is determined mostly by the circuit configuration. The resonant frequency for the magnetometer-readout scenario needs to be adjusted so that it matches the intrinsic resonant frequency of the Larmor precession frequency of the target gas. The resonant frequencies of the two readout scenarios exhibit variations owing to disparities in the circuit configurations. In particular, the inductance and capacitance of the Helmholtz coil in the magnetometer readout scenario contribute to a higher resonant frequency compared to the alternative ADC readout scenario. Among the various electric parameters, the resistance of the induction coil is identified as most dominant. This resistance depends on the frequency , mainly due to the presence of the skin effect and the proximity effect [dowell1966effects]. However, the exact relation between the coil resistance and signal frequency cannot be analytically given due to the complexity of the coil structure. To investigate the frequency dependence of the coil resistance, in-situ measurements are conducted using an HIOKI IM3533-01 LCR meter. The magnitude is denoted as , and the phase angle is denoted as . The and have correlation with the inductance , capacitance , and resistance of the coil, which can be expressed as:
@@
where is the frequency-independent direct resistance of the coil. The parameters and are empirical model parameters. A Nelder-Mead fitting algorithm is utilized to derive the , , and from the set of measured and . In a primary test, three induction coils with different radii and turn numbers are manufactured and tested. Their geometric parameters are given in Table 1. The derived for these induction coils (normalized to resistivity) are shown in Fig. 5. The waveforms of the induction electric current on the Helmholtz coil in magnetometer readout scenario and of electric current flowing into ADC in alternative direct readout scenario (the current in Fig. 3) are shown in Fig. 6, assuming the MM perpendicularly traverses the coil center of V1 with a speed of 10-5 light speed.